OpenStudy (anonymous):

combine like terms: 3xy-4/5 x^2y -2/9 xy+8/5 x^2y

OpenStudy (anonymous):

sorry these ones really confuse me :[

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

^2 means exponents 2

OpenStudy (cruffo):

"Like terms" means that the terms have to have the same variables and the same exponents. For example, \( 3x\) and \( 5x\) are like terms, but \(3x\) and \(3x^2\) are not.

OpenStudy (cruffo):

The same applies to multiple variables: \(3xy\) and \(5xy\) are like terms, but \(3xy\) and \(5xy^2\) are not

OpenStudy (cruffo):

\[\large 3xy-\frac{4}{5}x^2y-\frac{2}{9}xy+\frac{8}{5}x^2y\] \(\large 3xy\) and \(\large \frac{8}{5} xy\) are like terms and \(\large \frac{4}{5}x^2y\) and \(\large \frac{8}{5}x^2y\) are like terms

OpenStudy (cruffo):

sorry \(\large 3xy\) and \(\large \frac{2}{9} xy\) are like terms

OpenStudy (cruffo):

So what should happen when you combine the xy like terms? \[\large \color{red}{3xy}-\frac{4}{5}x^2y\color{red}{-\frac{2}{9}xy}+\frac{8}{5}x^2y\]